Present age of the father is 3 times that of his sons. The sum of that ages 10 years ago is as much as their present age of the father. Find the ages

Answer by checkley79(3341)

F=3X
S=X
3X-10+X-10=3X
3X-3X+X=10+10
X=20 AGE OF THE SON NOW.
3*20=60 AGE OF THE FATHER NOW.
PROOF:
60-10+20-10=60
80-20=60
60=60